7x-3=6x im not sure how to work this problem can you show me a break down thanks
Answer by Alan3354(69443)

7x-3=6x im not sure how to work this problem can you show me a break down thanks
------------------
You didn't say what you wanted, or needed, to do, but I'll assume it's to solve for x. Knowing what you want to do is step 1. If you don't know what you want to do, you'll never know when you've done it.
-----------------
Solve for x.
7x-3=6x
Solving for x means get x on one side of the equal sign by itself. That's the object of the game.
7x-3=6x
Subtract 6x from both sides.
x-3 = 0
Progress, right?
Now add 3 to both sides.
x = 3
We're done, take a break.